Someone invested $12,000.00 for her education. The person invests a portion in a bank certificate of deposit which earns 4% and the remainder in a savings bond that earns 7.5%. If the total interest earned after one year is $720.00, how much money was invested in the CD account?
Answer by checkley71(8403)

.04X+.075(12,000-X)=720
.04X+900-.075X=720
-.035X=720-900
-.035X=-180
X=-180/-.035
X=5,142.86 INVESTED @ 4%.
12,000-5,142.86=6,857.14
PROOF
.04*5,142.86+.075*6,857.14=720
205.71+514.29=720
720=720